OpenStudy (anonymous):

show ((sec2θ-1)/(2sec2θsinθ))=sinθ

OpenStudy (anonymous):

anyone can guide me for the first step?

OpenStudy (saifoo.khan):

is it sec^2 θ or sec2θ?

OpenStudy (anonymous):

no power

OpenStudy (saifoo.khan):

@TuringTest @UnkleRhaukus

OpenStudy (unklerhaukus):

\[\frac{\sec(2\theta)-1}{2\sec(2\theta)\sin(\theta)}\] \[=\frac{\frac1{\cos(2\theta)}-1}{2\frac1{\cos(2\theta)}\sin(\theta)}\] now multiply terms in the numerator And in the denominator by \(\cos(2\theta)\) and then use a half angle formula in the numerator

OpenStudy (anonymous):

Any problem you are facing @mathnoob1

OpenStudy (anonymous):

i having problem seeing the next step after UnkleRhaukus 1st step

OpenStudy (anonymous):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

\[=\frac{\frac1{\cos(2\theta)}-1}{2\frac1{\cos(2\theta)}\sin(\theta)} \times \frac{ \cos (2 \theta) }{ \cos (2 \theta) }\]\[=\frac{1-\cos(2 \theta)}{2\sin(\theta)}\]\[\frac{ 1-\cos(2 \theta) }{ 2 }=\sin^2\theta\]You can finish from here

OpenStudy (anonymous):

hmm i still lost from the 2nd step to the 3rd step..

OpenStudy (unklerhaukus):

\[=\frac{1-\cos(2 \theta)}{2\sin(\theta)}\] \[=\frac{1-\cos(2 \theta)}{2}\times\frac1{\sin(\theta)}\] this is the half angle formula \[\boxed{\frac{ 1-\cos(2 \theta) }{ 2 }=\sin^2\theta}\]